In šŸ”ŗ ABC, 2ca.sin(A-B+C)/2 is equal to... (a) aĀ²+bĀ²-cĀ² (b) cĀ²+aĀ²-bĀ²(c) bĀ²-cĀ²-aĀ² (D) cĀ²-aĀ²-bĀ²

Grade:11

1 Answers

Arun
6 years ago
A+B+C=Pi

so A+C = Pi - B

so here it will be :Ā  2ac.sin(Pi/2 -B) = 2ac.cosB

by cosine rule, we have: c2+a2-b2Ā = 2ac cos B

optionĀ B